OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. Show your work. f(x) =

OpenStudy (anonymous):

@dan815 @abb0t

OpenStudy (anonymous):

@DullJackel09 @TheRaggedyDoctor

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@poopsiedoodle @pooja195

OpenStudy (anonymous):

plz help me id understand this trickery at all

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

okay pick a rational function and tell me

OpenStudy (anonymous):

what is a rational function lol

OpenStudy (anonymous):

haha

OpenStudy (freckles):

Consider where the fraction is not defined. When you have a fraction, what is that famous number you cannot divide by?

OpenStudy (anonymous):

0 lol i think or im stupid lol

OpenStudy (freckles):

You aren't stupid. And that is an awesome answer because it is true. So when is your fraction's bottom zero? For what value of x, do you have x-1 is zero?

OpenStudy (anonymous):

good question idk how to answer them

OpenStudy (freckles):

what number can you take away 1 from and have zero?

OpenStudy (freckles):

is 5-1 zero? is 2-1 zero?

OpenStudy (anonymous):

no 1-1 is 0

OpenStudy (freckles):

yes x-1=0 has solution x=1 you could have also just added 1 on both sides

OpenStudy (anonymous):

true ok i get it somewhat now

OpenStudy (anonymous):

ok so if i was to answer this question then how would i phrase this

OpenStudy (freckles):

x=1 is your vertical asymptote though if there was a factor of x-1 on top then it could possibly be a hole (but you do not have that here) that is f(x)=(x-1)/(x-1) has a hole at x=1 (not a vertical asymptote) that is f(x)=x(x-1)/(x-1) also has a hole at x=1 but g(x)=(x-1)/(x-1)^2 has a vertical asympote at x=1 since it has more (x-1)'s on bottom then on top

OpenStudy (anonymous):

i hate math so bad lol good thing this is my lsast math class lol thnx for your help

OpenStudy (freckles):

np

OpenStudy (anonymous):

thank you to dan